OpenStudy (bakonloverk):

the area of a triangle is 64square meters and the base is one-half times the height. find the triangles height

OpenStudy (anonymous):

First try writing out two equations, one that describes the area of a triangle, and then one that describes the relation between this triangle's base and height, as given.

OpenStudy (bakonloverk):

kk so area is base* height *.5= area

OpenStudy (anonymous):

Good, b*h*1/2 = a. How about another for the relation between the base and height of this triangle?

OpenStudy (anonymous):

Let h = height, then b = (1/2)h A = (1/2)bh 64 = (1/2)(1/2)h*h 64*4 = h^2 256 = h^2 h = 16

OpenStudy (bakonloverk):

wait so its 16?

OpenStudy (anonymous):

here you know the formula for area correct. well plug in what you know.\[64=\frac{ 1 }{ 2 }(\frac{ x }{ 2 })x\]

OpenStudy (bakonloverk):

but then wht?

OpenStudy (anonymous):

then solve, \[64(\frac{ 2 }{ 1 })=(\frac{ x }{ 2 })(\frac{ x }{ 1 })\]

OpenStudy (bakonloverk):

i got 128 but idk

OpenStudy (anonymous):

\[128=\frac{ x ^{2} }{ 2 }\] you combine the right side

OpenStudy (bakonloverk):

so multiply 128 but 2 and then find the square root?

OpenStudy (anonymous):

yep you got it.